### v1

# Efficient, Accurate and Stable Gradients for Neural ODEs

Abstract
Neural ODEs are a recently developed model class that combine the strong model priors of differential equations with the high-capacity function approximation of neural networks. One advantage of Neural ODEs is the potential for memory-efficient training via the continuous adjoint method. However, memory-efficient training comes at the cost of approximate gradients. Therefore, in practice, gradients are often obtained by simply backpropagating through the internal operations of the forward ODE solve - incurring high memory cost.
Interestingly, it is possible to construct algebraically reversible ODE solvers that allow for both exact gradients and the memory-efficiency of the continuous adjoint method. Unfortunately, current reversible solvers are low-order and suffer from poor numerical stability. The use of these methods in practice is therefore limited.
In this work, we present a class of algebraically reversible solvers that are both high-order and numerically stable. Moreover, any explicit numerical scheme can be made reversible by our method. This construction naturally extends to numerical schemes for Neural CDEs and SDEs.
1 Introduction
Neural ODEs
Neural Ordinary Differential Equations (ODEs) introduce a model class where the vector field of an ODE is parameterised as a neural network [1]. The continuous time model admitted by this construction has interesting benefits over discrete time models, such as adaptive computation and memory-efficient training. Additionally, Neural ODEs and their extensions to Neural CDEs and SDEs, have been found to outperform discrete models in many applications including scientific, generative and time-series modelling [2, 3, 4, 5, 6, 7, 8]. Indeed, many discrete models are discretisations of differential equations - such as residual networks [9], recurrent networks [10] and momentum residual networks [11].
Training
Training Neural ODEs proceeds by two main approaches: standard backpropagation through the internal numerical operations (discretise-then-optimise) or via the continuous adjoint method (optimise-then-discretise) [12].
Discretise-then-optimise is the generally preferred training method as the parameter gradients obtained are exact. That is, the gradients correspond exactly to the numerical model being trained. However, it is required that we store the forward computation graph in memory; and this can quickly become prohibitively expensive.
Alternatively, optimise-then-discretise does not require that we store the forward computation graph, only the terminal ODE state. The memory-efficiency of training is therefore significantly improved. However, gradients obtained by optimise-then-discretise are a numerical approximation to the true gradients. This approximation can cause problems during training and models trained via optimise-then-discretise may fail to converge.
Algebraically reversible solvers
Interestingly, it is possible to achieve both the memory efficiency of optimise-then-discretise and obtain the exact gradients of discretise-then-optimise. This is realised by constructing an algebraically reversible numerical solver, whereby the solver state at step $n$ can be reconstructed in closed form from the solver state at step $n+1$.
There have been two previously proposed algebraically reversible solvers: Zhuang et al. [13] introduce MALI that is a second-order reversible ODE solver based on the asynchronous leap-frog method; and, Kidger et al. [14] devise a reversible Heun method that is a second-order ODE solver and $0.5$ strong order SDE solver.
However, both reversible solvers are low-order and suffer from poor numerical stability. The use of these methods in practice is therefore limited.
1.1 Contributions
In this work we address both the low-order convergence and stability issues of previous reversible solvers for Neural ODEs.
We present a general class of algebraically reversible solvers that allows any explicit numerical solver to be made reversible. This class of reversible solvers produce exact, memory-efficient gradients and are:
  1. 1.
    high-order (see Theorem 3.1),
  2. 2.
    numerically stable (see Theorem 3.2),
  3. 3.
    and naturally extend to Neural CDEs and SDEs.
2 Background
2.1 Neural ODEs
A Neural ODE is defined as [1]
$\frac{\mathrm{d}y}{\mathrm{d}t}(t)=f_{\theta}(t,y(t)),\hskip 22.76219pty(0)=y_{0},$ (1)
where $y(t)\in\mathbb{R}^{d}$ is the solution to (1) and $f_{\theta}:\mathbb{R}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a neural network parameterised by $\theta\in\mathbb{R}^{m}$. The initial condition is given by $y_{0}\in\mathbb{R}^{d}$.
For example, to perform regression with (1) over a dataset of input-output pairs $\mathcal{D}=\{(x_{i},y_{i})\}_{0\leq i\leq n}$, we let $y(0)=x_{i}$ and solve the ODE forward in time to compute $y(T)\approx y_{i}$ for some terminal time $T$. The parameters $\theta$ are learnt to minimise a loss function between $y(T)$ and $y_{i}$, over the dataset.
2.2 Training
Training Neural ODEs requires backpropagating through the forward solve and therefore corresponds to a backwards-in-time operation. The two main methods for achieving this are to backpropagate the numerical forward-solve directly (discretise-then-optimise) or to solve the continuous adjoint equations (optimise-then-discretise) [12].
Out setup is as follows: Let $L(y(T)):\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a scalar-valued loss function dependent on the terminal state $y(T)$. We wish to obtain the gradient of $L(y(T))$ with respect to the vector field parameters $\theta$, given by $\mathrm{d}L/\mathrm{d}\theta$.
2.2.1 Discretise-then-Optimise
Provided our ODE solver is written in a programming language that supports automatic differentiation, it is straightforward to backpropagate through the internal operations of the solver. This approach is called discretise-then-optimise.
Exact gradients
The gradients $\mathrm{d}L/\mathrm{d}\theta$ obtained by discretise-then-optimise will be exact in the sense that the exact numerical values used on the forward solve will be used to calculate the gradient.
High memory cost
However, the computation graph of the forward solve is required to be stored in memory. The memory cost of this graph scales as $\mathcal{O}(NM)$, where $N$ is the number of solver steps taken and $M$ is the cost of one step.
Checkpointing
The memory cost may be reduced by saving the computation graph at a set of checkpoints and reconstructing the graph from these points [15]. The memory cost is $\mathcal{O}(CM)$, where $C$ is the number of checkpoints. However, the number of checkpoints is typically a function of the number of solver steps $C(N)$ and therefore the memory cost still scales with $N$.
Additionally, recomputing the forward solve from checkpointed values, while maintaining the $\mathcal{O}(CM)$ memory cost introduces a significant computation-time cost. Specifically, the computation-time cost of backpropagating $L$ steps from one checkpointed value is quadratic, $\mathcal{O}(L^{2})$ [16].
2.2.2 Optimise-then-Discretise
A memory-efficient alternative for backpropagating through the forward solve is via the continuous adjoint method. An implicit expression for the gradient $\mathrm{d}L/\mathrm{d}\theta$ can be obtained by appropriately differentiating the Neural ODE (1). This approach is called optimise-then-discretise.
The implicit expression for $\mathrm{d}L/\mathrm{d}\theta$ is in the form of another ODE, called the adjoint ODE. For the Neural ODE in (1), it can be shown that the adjoints $a_{\theta}(t)=\mathrm{d}L/\mathrm{d}\theta(t)$ and $a_{y}(t)=\mathrm{d}L/\mathrm{d}y(t)$ satisfy
$\frac{\mathrm{d}a_{y}}{\mathrm{d}t}(t)$ $=-a_{y}(t)^{\mathrm{T}}\frac{\partial f_{\theta}}{\partial y}(t,y(t)),\hskip 22.76219pta_{y}(T)=\frac{\mathrm{d}L}{\mathrm{d}y(T)}$ (2)
$\frac{\mathrm{d}a_{\theta}}{\mathrm{d}t}(t)$ $=-a_{y}(t)^{\mathrm{T}}\frac{\partial f_{\theta}}{\partial\theta}(t,y(t)),\hskip 22.76219pta_{\theta}(T)=0.$
The system of coupled ODEs in (2) is numerically solved backwards in time to obtain $a_{\theta}(0)=\mathrm{d}L/\mathrm{d}\theta$ [17].
Low memory cost
The computation graph of the forward solve is no longer required, only the terminal state of the solver is necessary. The memory cost of optimise-then-discretise is therefore $\mathcal{O}(M)$, where $M$ is the cost of one solver step.
Approximate gradients
However, solving the system in (2) requires $y(t)$ which must be numerically re-approximated on the backward solve. The values of $y(t)$ obtained on the backward solve are different to those produced on the forward solve. The gradients calculated are therefore a numerical approximation of the true gradients and this can affect model training.
3 Reversible Solvers
We now introduce a class of reversible solvers, where any explicit numerical solver can be made reversible. The algebraic reversibility allows one to dynamically recompute the forward solve in closed form during backpropagation, thereby obtaining exact gradients and requiring only the terminal state of the solver be stored in memory.
Forward Solve
Suppose we have an explicit numerical ODE solver, $\Psi_{h}(t,y):\mathbb{R}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, such that a numerical step is given by $y_{n+1}=y_{n}+\Psi_{h}(t,y)$, where $h$ is the step size. Then we construct a reversible numerical solution $\{y_{n},z_{n}\}_{n\geq 0}$ by
$y_{n+1}$ $=\lambda y_{n}+(1-\lambda)z_{n}+\Psi_{h}(t_{n},z_{n}),$ (3)
$z_{n+1}$ $=z_{n}-\Psi_{-h}(t_{n+1},y_{n+1}),$
where $\lambda\in(0,1]$ is a coupling parameter. The initial state is given by $z(0)=y(0)=y_{0}$ and $t_{n}=nh$ for each $n\geq 0$.
Backward Solve
The numerical scheme in (3) is algebraically reversible, with the reverse scheme given by
$z_{n}$ $=z_{n+1}+\Psi_{-h}(t_{n+1},y_{n+1}),$ (4)
$y_{n}$ $=\lambda^{-1}y_{n+1}+\left(1-\lambda^{-1}\right)z_{n}-\lambda^{-1}\Psi_{h}(t_{n},z_{n}).$
Coupling parameter
The choice of coupling parameter $\lambda$ is key to the reversible method. Stability is increased on the forward solve by decreasing $\lambda$ at the cost of stability on the backward solve, due to the $\lambda^{-1}$ multiplier. It is therefore important that $\lambda$ remains close to $1$ and in practice we find $\lambda\in[0.99,0.999]$ to be good choices. See Section 3.2 for more details.
Convergence
The solver in (3) is not only explicit and reversible, but provided $\Psi$ is a method with $k$-th order convergence then the reversible scheme will also have $k$-th order convergence. This is shown by the following theorem. For this reason we refer to the reversible scheme with $\Psi$ from solver $X$ as ‘Reversible $X$’.
Theorem 3.1.
For a fixed time-horizon $T>0$, we consider the Neural ODE in (1) over $[0,T]$. Let $T=Nh$ where $N>0$ denotes the number of steps and $h>0$ is the step size.
Let $\Psi$ be an order $k$ ODE solver satisfying the Lipschitz condition (see A.2) and consider the reversible solution $\{y_{n},z_{n}\}_{n\geq 0}$, given by
$y_{n+1}$ $=\lambda y_{n}+(1-\lambda)z_{n}+\Psi_{h}(t_{n},z_{n}),$
$z_{n+1}$ $=z_{n}-\Psi_{-h}(t_{n+1},y_{n+1}),$
where $\lambda\in(0,1]$ is a coupling parameter. The initial state is given by $z(0)=y(0)=y_{0}$ and $t_{n}=nh$ for each $n\geq 0$. Then there exist constants $h_{\text{max}},C>0$ such that, for $h\in[0,h_{\text{max}}]$,
$||y_{n}-y(t_{n})||\leq Ch^{k}.$ (5)
Proof.
The proof of Theorem 3.1 is given in Appendix A. ∎
3.1 Memory-efficient backpropagation
Due to the algebraic reversibility of the solver, we can perform memory-efficient backpropagation through the forward solve. The key point is that the gradients obtained are exactly the discretise-then-optimise gradients and require only $\mathcal{O}(M)$ memory, where $M$ is the memory cost of one solver step.
The backpropagation algorithm through one solver step is given in Algorithm 1 and follows the construction of reverse-mode automatic differentiation [18]. Our setup is as follows.
Let $y_{n},z_{n}\in\mathbb{R}^{d}$ and define a scalar-valued loss function $L(y_{N}):\mathbb{R}^{d}\rightarrow\mathbb{R}$, where $y_{N}\approx y(T)$ is the terminal value of the numerical solve. The network parameters are given by $\theta\in\mathbb{R}^{m}$.
We define the adjoint quantities $\bar{v}=\partial L(y_{N})/\partial v$ for ease of presentation.
Algorithm 1 Memory-efficient Backpropagation
$t_{n+1},y_{n+1},z_{n+1},\bar{y}_{n+1},\bar{z}_{n+1},\bar{\theta}$
# Reverse step
$z_{n}=z_{n+1}+\Psi_{h}(t_{n+1},y_{n+1})$
$y_{n}=\lambda^{-1}y_{n+1}+\left(1-\lambda^{-1}\right)z_{n}-\lambda^{-1}\Psi_{h}(t_{n},z_{n})$
# Gradients
$\bar{y}_{n+1}\leftarrow\bar{y}_{n+1}-\bar{z}_{n+1}\displaystyle\frac{\partial\Psi_{-h}(t_{n+1},y_{n+1})}{\partial y_{n+1}}$
$\bar{y}_{n}=\lambda\bar{y}_{n+1}$
$\bar{z}_{n}=\bar{z}_{n+1}+(1-\lambda)\bar{y}_{n+1}+\bar{y}_{n+1}\displaystyle\frac{\partial\Psi_{h}(t_{n},z_{n})}{\partial z_{n}}$
$\bar{\theta}\leftarrow\bar{\theta}-\bar{z}_{n+1}\displaystyle\frac{\partial\Psi_{-h}(t_{n+1},y_{n+1})}{\partial\theta}+\bar{y}_{n+1}\displaystyle\frac{\Psi_{h}(t_{n},z_{n})}{\partial\theta}$
$t_{n},y_{n},z_{n},\bar{y}_{n},\bar{z}_{n},\bar{\theta}$
Implementation
The multiplication of the adjoint quantities $\bar{v}$ with derivatives of $\Psi$ in Algorithm 1 can be efficiently computed using vector-Jacobian products. This requires no knowledge on the structure of $\Psi$ a-priori and no local forward evaluations as $\Psi$ is evaluated on the backward solve.
3.2 Stability
One key ingredient missing from previous reversible solvers is numerical stability. Here we show that the reversible scheme introduced above in (3) has a non-zero linear stability region for $\Psi$ given by any explicit Runge-Kutta (RK) method. This is shown by Theorem 3.2.
We prove stability, as is customary, for the real-valued linear test problem [19].
Definition 3.1 (Linear Stability).
Let $y\in\mathbb{R}^{d}$ and consider the ODE
$\frac{\mathrm{d}y}{\mathrm{d}t}=\alpha y,\hskip 22.76219pty(0)=y_{0},$ (6)
where $\alpha<0$ is a negative real constant and $y_{0}\in\mathbb{R}^{d}$ is a non-zero initial condition. A numerical solution $\{y_{n}\}_{n\geq 0}$ to (6) is linearly stable if the spectral radius (maximum eigenvalue) $\rho(T)<1$, where $T\in\mathbb{R}^{d\times d}$ is given by
$y_{n+1}=Ty_{n}.$ (7)
Transfer functions
It is well known that any explicit Runge-Kutta method applied to (6) admits a transfer function $T(h\alpha)$ such that $y_{n+1}=T(h\alpha)y_{n}$, where $h$ is the step size [20]. Equivalently, we may express this in terms of the propagating function $\Psi_{h}(t_{n},y_{n})=R(h\alpha)y_{n}$, where $R(h\alpha)=T(h\alpha)-1$.
Theorem 3.2.
Let $\Psi$ be given by an explicit Runge-Kutta solver. Then the reversible numerical solution $\{y_{n},z_{n}\}_{n\geq 0}$ given by
$y_{n+1}$ $=\lambda y_{n}+(1-\lambda)z_{n}+\Psi_{h}(t_{n},z_{n}),$
$z_{n+1}$ $=z_{n}-\Psi_{-h}(t_{n+1},y_{n+1}),$
is linearly stable iff
$|\Gamma|<1+\lambda,$ (8)
where
$\Gamma=1+\lambda-(1-\lambda)R(-h\alpha)-R(-h\alpha)R(h\alpha).$ (9)
Proof.
The proof of Theorem 3.2 is given in Appendix B. ∎
Comparison to previous methods
It should be noted that the second-order reversible solver introduced in [13] and the reversible Heun method introduced in [14] are nowhere linearly stable, for any step size $h$. In contrast, Theorem 3.2 shows that for coupling parameter $\lambda\in(0,1)$ there exists a step size $h$ such that the reversible scheme presented here is linearly stable.
Examples
[…]
We see that the stability region is decreased in comparison to the original solver but is non-zero in the left half complex plane. This is in stark contrast to previous reversible methods, where the stability region is confined to the imaginary axis $[-i,i]$. In practice, we find that the non-zero A-stability region ensures that the reversible method is numerically stable for relatively large step sizes $h$.
[…]
4 Experiments
The performance of the reversible solver was validated in three experiments.
Toy problem
The first is a toy regression problem to isolate the measurement of gradient error and memory usage compared across the reversible method, optimise-then-discretise and discretise-then-optimise.
Continuous Normalising Flow
In the second experiment, we apply the reversible solver to train a Continuous Normalising Flow (CNF) [21], an application where memory usage of discretise-then-optimise can become prohibitively high.
Neural CDE
The third experiment trains a Neural CDE, demonstrating the applicability of the solver to non-ODE based models.
4.1 Toy problem
We begin by measuring the gradient error and memory usage of the reversible method. The key validation criteria are that the reversible solver recovers exactly the gradients used by discretise-then-optimise in constant memory with respect to the depth of the computation graph.
The Neural ODE was trained on a polynomial regression task with the Reversible Midpoint method. The optimise-then-discretise and disretise-then-optimise training methods also used the (non-reversible) midpoint scheme. Precise experimental details can be found in Appendix C.
The resulting gradient error and memory usage are shown in Figure 2.
Gradient error
In Figure 2(a), the $L^{1}$ gradient error of the Reversible Midpoint method and optimise-then-discretise are compared to the discretise-then-optimise ground truth.
We see that the gradient error of the reversible method is constant with integration time and the value of the error is zero (within a negligible reconstruction error). In comparison, the gradient error of optimise-then-discretise increases linearly with integration time and the value of the error becomes large.
Memory usage
In Figure 2(b), the memory usage of the Reversible Midpoint method is compared to the memory usage of discretise-then-optimise.
We see that the memory usage is constant with respect to integration time. This is expected as the method only requires the final solver state to be stored in memory, which is independent of integration time. In comparison, the memory usage of discretise-then-optimise increases linearly with integration time in line with the size of the forward computation graph.
Figure 2: (a) $L^{1}$ gradient error of the reversible method and optimise-then-discretise compared to discretise-then-optimise ground truth. (b) Memory usage of the reversible method during backpropagation compared to discretise-then-optimise.
4.2 Continuous Normalising Flow
The second experiment we perform is to train a Continuous Normalising Flow (CNF) [21]. CNFs are used to model arbitrary probability distributions. This is achieved by learning the push-forward map from a simpler known distribution (usually Gaussian).
Advantages of Reversible Solver
CNFs are a great application for reversible solvers as commonly CNFs are used to model a distribution over high dimensional objects (such as images). The memory cost of storing the objects state during integration, for discretise-then-optimise backpropagation, can often become prohibitively high.
Memory-efficient backpropagation
We also introduce a further memory-efficiency improvement for training CNFs. The improvement is a simple algorithm tweak that is yet to be exploited by previous methods.
4.2.1 Memory-efficient backpropagation
Training a CNF proceeds by minimising the negative log-likelihood of the data. Consider the Neural ODE in (1) and an individual data sample $y\in\mathbb{R}^{d}$. The log-likelihood is given by [12, 21]
$\log p_{\theta}(T,y)=\log p_{\theta}(0,y(0))-\int_{0}^{T}\sum_{k=1}^{d}\frac{\partial f_{\theta,k}}{\partial y_{k}}(t,y(t))\mathrm{d}t,$ (10)
where $p_{\theta}:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that $p_{\theta}(t,\cdot)$ is the density of $y(t)$ for each time $t\in[0,T]$. The subscript $k$ indicates the $k$-th element of a vector.
Obtaining (10) requires that we solve the Neural ODE backwards in time to get $y(0)$. Typically, the integral term is calculated alongside the Neural ODE solve, with the vector field given by the integrand. That is, the integral is treated as an ODE and solved accordingly.
However, this decision introduces computational redundancy during backpropagation as the calculation of the integral term is intrinsically algebraically reversible (unlike standard numerical ODE solvers). Storing the integral computation, as is done in discretise-then-optimise, is unnecessarily memory-intensive; and numerically re-approximating it, as is done in optimise-then-discretise, introduces unnecessary error.
In Algorithm 2 we present the memory-efficient backpropagation algorithm through one reversible solver step, where the intrinsic reversibility of the integral calculation is utilised. Note that as (10) is solved backwards in time, backpropagation corresponds to a forward-in-time operation. The setup remains identical to Algorithm 1 with an additional state for the integral term. Let $I_{n}\in\mathbb{R}$ be the integral state at step $n$ with solver step function $\Phi:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$.
Algorithm 2 Continuous Normalising Flow Backpropagation
$t_{n},y_{n},z_{n},I_{n},\bar{y}_{n},\bar{z}_{n},\bar{I}_{n},\bar{\theta}$
# Forward step
$y_{n+1}=\lambda y_{n}+(1-\lambda)z_{n}+\Psi_{h}(t_{n},z_{n})$
$z_{n+1}=z_{n}-\Psi_{-h}(t_{n+1},y_{n+1})$
$I_{n+1}=I_{n}-\Phi_{-h}(t_{n+1},y_{n+1})$
# Gradients
$\bar{z}_{n}\leftarrow\bar{z}_{n}+(1-\lambda^{-1})\bar{y}_{n}-\lambda^{-1}\bar{y}_{n}\displaystyle\frac{\partial\Psi_{h}(t_{n},z_{n})}{\partial z_{n}}$
$\bar{z}_{n+1}=\bar{z}_{n}$
$\bar{y}_{n+1}=\bar{z}_{n}\displaystyle\frac{\Psi_{-h}(t_{n+1},y_{n+1})}{\partial y_{n+1}}+\lambda^{-1}\bar{y}_{n}+\bar{I}_{n}\displaystyle\frac{\partial\Phi_{-h}(t_{n+1},y_{n+1})}{\partial y_{n+1}}$
$\bar{I}_{n+1}=\bar{I}_{n}$
$\bar{\theta}\leftarrow\bar{\theta}+\bar{z}_{n}\displaystyle\frac{\partial\Psi_{-h}(t_{n+1},y_{n+1})}{\partial\theta}-\lambda^{-1}\bar{y}_{n}\displaystyle\frac{\partial\Psi_{h}(t_{n},z_{n})}{\partial\theta}+\bar{I}_{n}\displaystyle\frac{\partial\Phi_{-h}(t_{n+1},y_{n+1})}{\partial\theta}$
$t_{n+1},y_{n+1},z_{n+1},I_{n+1},\bar{y}_{n+1},\bar{z}_{n+1},\bar{I}_{n+1},\bar{\theta}$
4.2.2 Results
The reversible solver method was used to train a CNF model on the two moons dataset [22]. Additionally, for comparison, the CNF model was trained using discretise-then-optimise. The resulting negative log-likelihoods, $-\mathbb{E}\left[\log p_{\theta}\right]$, and memory usages are shown in Table 1, for a selection of numerical solvers. Further experimental details can be found in Appendix C.
Log-likelihood
The negative log-likelihood obtained by the reversible method is consistently lower than or comparable to that obtained by discretise-then-optimise. This result is consistent across a selection of numerical solvers.
Memory usage
The reversible method significantly outperforms discretise-then-optimise on memory usage. The improvement is most considerable for higher-order methods, where the number of intermediate calculations becomes large and discretise-then-optimise must store these intermediate values at each time step.
It should be noted that there exists a base memory cost as a result of the just-in-time compiled JAX functions. This base cost is comparable across both methods.
Table 1: Continuous Normalising Flow on the two moons dataset; trained via the reversible method and compared to discretise-then-optimise. Mean $\pm$ standard deviation averaged over $10$ runs.
Solver $-\mathbb{E}\left[\log p_{\theta}\right]$ Memory Usage (GB)
Reversible Discretise-then-Optimise Reversible Discretise-then-Optimise
Midpoint $\mathbf{0.888\pm 0.019}$ $0.891\pm 0.017$ $\mathbf{0.563\pm 0.002}$ $3.922\pm 0.001$
RK4 $\mathbf{0.890\pm 0.016}$ $0.890\pm 0.018$ $\mathbf{0.647\pm 0.002}$ $7.467\pm 0.001$
Dopri5 $\mathbf{0.890\pm 0.017}$ $0.891\pm 0.018$ $\mathbf{0.704\pm 0.003}$ $12.788\pm 0.002$
4.3 Neural CDE
Neural CDEs are an extension of Neural ODEs to time-series modelling [5]. A limitation of Neural ODEs is that the dynamics are completely determined by the initial condition. Neural CDEs fix this problem by allowing the ODE state to be continuously controlled by an input path.
CharacterTrajectories dataset
We apply the Neural CDE model to the CharacterTrajectories dataset [23]. The data consists of the $(x,y)$ velocity trajectory and pen tip force while drawing a Latin alphabet character. The reversible and discretise-then-optimise methods were used to train the Neural CDE model for a selection of numerical solvers. The accuracies and memory usages of the two methods are shown in Table 2.
Further experimental details can be found in Appendix C.
Accuracy
The accuracy obtained by the reversible solver is consistently higher than or comparable to the accuracy obtained by discretise-then-optimise. This effect is maintained over a selection of numerical solvers.
Memory Usage
The memory usage of the reversible method is significantly decreased in comparison to discretise-then-optimise. For higher order solvers, this effect is most pronounced as discretise-then-optimise must store an increased number of intermediate calculations.
Table 2: Neural CDE on CharacterTrajectories dataset; trained via the reversible method and compared to discretise-then-optimise. Mean $\pm$ standard deviation averaged over $10$ runs.
[…]
Reversible Discretise-then-Optimise Reversible Discretise-then-Optimise
Midpoint $78.4\pm 5.5$ $\mathbf{78.9\pm 6.7}$ $\mathbf{0.434\pm 0.001}$ $1.09\pm 0.001$
RK4 $\mathbf{79.0\pm 5.9}$ $76.4\pm 5.4$ $\mathbf{0.468\pm 0.001}$ $1.86\pm 0.001$
Dopri5 $\mathbf{80.1\pm 6.9}$ $77.9\pm 6.7$ $\mathbf{0.523\pm 0.002}$ $3.01\pm 0.002$
5 Discussion
5.1 Implementation
JAX
An implementation of the reversible solver method was written in JAX [24]. The code is available at https://github.com/sammccallum/reversible. We are also currently working to implement the reversible solver method in the popular JAX differential equation solving library Diffrax [12].
5.2 Performance tricks
Coupling parameters and step sizes
The linear stability region for an explicit numerical method solving forwards in time is the inverse of the stability region when solving backwards in time. Therefore, overall stability may be improved by straddling the edge of the stability region (known from stability analysis, Section 3.2). For high-order reversible solvers, we found that this was not required and numerical stability was easily obtained with any sensible choice of $\lambda$ and $h$.
Weight decay
It is well known in the Neural ODE literature that adding weight decay to the vector field parameters can sufficiently constrain the problem to be non-stiff [21]. This result was also verified here; numerical stability improved with weight decay.
5.3 Limitations
Computation speed
Due to the reconstruction of the forward numerical solve during backpropagation, the reversible method is slower than discretise-then-optimise. At each time-step, the reversible method requires one additional solver step evaluation on the forward solve and two additional solver step evaluations on the backward solve, in comparison to discretise-then-optimise.
Stability
[…]
5.4 Future work
Adaptive step sizes
The reversible method may lend itself to the use of adaptive step sizes due to the additional solver state. A calculation of the numerical error can be directly obtained by comparing the states ($y_{n},z_{n}$) without requiring an embedded numerical method.
Partial Differential Equations
The finite difference approximation to PDEs results in a set of coupled ODEs. The number of ODEs is equal to the size of the discretisation grid and the forward computation graph can therefore incur high memory cost [25]. Application of reversible solvers to discretised PDEs would significantly decrease this memory cost.
Implicit Models
Neural ODEs have recently been situated in the larger model class of Implicit models. For example, Deep Equilibrium Models are an example of an implicit model where the solution is specified as the fixed/equilibrium point of some system [26]. Implicit models require a numerical scheme to determine the solution. If one can devise a numerical scheme that is algebraically reversible, then the memory-efficiency and gradient accuracy of reversible differential equation solvers may be extended to other implicit models.
6 Conclusion
A general class of reversible solvers was introduced that provide memory-efficient and exact gradient calculation for training Neural ODE models. The reversible solvers presented are high-order, numerically stable and naturally extend to Neural CDEs and SDEs. The performance of the reversible method was shown to significantly outperform optimise-then-discretise and discretise-then-optimise on gradient accuracy and memory usage, respectively.
Appendix
Appendix A Proof of convergence (Theorem 3.1)
We will consider the following ODE,
$\frac{dy}{dt}$ $=f(y),$ (11)
$y(0)$ $=y_{0},$
where the state $(t,y)$ has been concatenated to $y\in\mathbb{R}^{n}$ without loss of generality. The vector field is given by the Lipschitz function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and the initial condition is $y_{0}\in\mathbb{R}^{n}$.
[…]
Definition A.1.
For $h_{\max}>0$, consider the map $\Psi$ given by
$\Psi:[-h_{\max},h_{\max}]\times\mathbb{R}^{n}$ $\rightarrow\mathbb{R}^{n},$
$(h,x)$ $\mapsto\Psi_{h}(x).$
We say that $\Psi$ is an ODE solver with order $\alpha>0\,$ if there exists a constant $C_{1}>0$ such that
$\big{\|}x(h)-\big{(}x+\Psi_{h}(x)\big{)}\big{\|}$ $\leq C_{1}|h|^{\alpha+1},$ (12)
for $h\in[-h_{\max},h_{\max}]\,$ where $x(h)$ denotes the solution $y(t)$ at time $t=|h|$ of the ODE:
$y^{\prime}$ $=\mathrm{sgn}(h)\cdot f(y),$
$y(0)$ $=x.$
Definition A.2.
Let $\Psi$ denote an order $\alpha$ solver for an ODE governed by a Lipschitz vector field $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. We say that $\Psi$ is an ODE solver satisfying the Lipschitz condition if there exists a constant $C_{2}>0$ such that
$\big{\|}\Psi_{h}(x)-\Psi_{h}(y)\big{\|}\leq C_{2}|h|\big{\|}x-y\big{\|},$ (13)
for all $x,y\in\mathbb{R}^{n}$ and $h\in[-h_{\max},h_{\max}]$. In particular, this may be written as
$\big{\|}\Psi_{h}\big{\|}_{\text{Lip-}1}\leq C_{2}|h|.$ (14)
We note that the Lipschitz condition defined in A.2 is true for all explicit Runge-Kutta methods; this result is shown in Section $6.1.4.2$ of [27].
Lemma A.1.
Let $\Psi$ denote an order $\alpha$ ODE solver satisfying the Lipschitz condition. Then there exists a constant $C_{3}>0$ such that
$\big{\|}\Psi_{h}(x+\Psi_{-h}(x))+\Psi_{-h}(x)\big{\|}$ $\leq C_{3}|h|^{\alpha+1},$
for $h\in[-h_{\max},h_{\max}]$.
Proof.
By Definition A.1, A.2 and the triangle inequality, we have
$\big{\|}\Psi_{h}(x+\Psi_{-h}(x))+\Psi_{-h}(x)\big{\|}$ $=\big{\|}\big{(}x(-h)+\Psi_{h}(x(-h))-x\big{)}+\big{(}x+\Psi_{-h}(x)-x(-h)\big{)}$
$\hskip 28.45274pt+\Psi_{h}(x+\Psi_{-h}(x))-\Psi_{h}(x(-h))\big{\|}$
$\leq\big{\|}x(-h)+\Psi_{h}(x(-h))-x\big{\|}+\big{\|}x+\Psi_{-h}(x)-x(-h)\big{\|}$
$\hskip 28.45274pt+\big{\|}\Psi_{h}(x+\Psi_{-h}(x))-\Psi_{h}(x(-h))\big{\|}$
$\leq 2C_{1}|h|^{\alpha+1}+C_{2}|h|\big{\|}x+\Psi_{-h}(x)-x(-h)\big{\|}$
$\leq\underbrace{C_{1}\big{(}2+C_{2}\,h_{\max}\big{)}}_{=:\,C_{3}}|h|^{\alpha+1}.$
Next, we bound the error between the reversible solver states in lemma A.2.
Lemma A.2.
For a fixed time horizon $T>0$, we consider the ODE (11) over the interval $[0,T]$. Suppose that $T=Nh$ were $N>0$ denote the number of steps and $h>0$ denotes the step size. Consider the reversible numerical solution $\{(Y_{k},Z_{k})\}_{0\leq k\leq N}$ with $Y_{0}=Z_{0}=y_{0}$ and, for $k\geq 0$,
$Y_{k+1}$ $:=\lambda Y_{k}+(1-\lambda)Z_{k}+\Psi_{h}(Z_{k}),$
$Z_{k+1}$ $:=Z_{k}-\Psi_{-h}(Y_{k+1}),$
where $\lambda\in(0,1]$ and $\Psi$ denotes an order $\alpha$ ODE solver satisfying the Lipschitz condition. Then there exists $h_{\max},C_{4}>0$ such that
$\|Y_{k}-Z_{k}\|\leq C_{4}h^{\alpha}\,,$ (15)
for $h\in(0,h_{\max}]$ and $0\leq k\leq N$.
Proof.
Letting $E_{k}^{YZ}:=\|Y_{k}-Z_{k}\|$, it follows from definition A.2 and lemma A.1 that
$E_{k+1}^{YZ}$ $=\|Y_{k+1}-Z_{k+1}\|$
$=\big{\|}\lambda\big{(}Y_{k}-Z_{k}\big{)}+\Psi_{h}(Z_{k})+\Psi_{-h}(Y_{k+1})\big{\|}$
$=\big{\|}\lambda\big{(}Y_{k}-Z_{k}\big{)}+\Psi_{h}(Z_{k})+\Psi_{-h}(Z_{k+1})+\Psi_{-h}(Y_{k+1})-\Psi_{-h}(Z_{k+1})\big{\|}$
$\leq\lambda\|Y_{k}-Z_{k}\|+\big{\|}\Psi_{h}(Z_{k})+\Psi_{-h}(Z_{k+1})\big{\|}+\big{\|}\Psi_{-h}(Y_{k+1})-\Psi_{-h}(Z_{k+1})\big{\|}$
$\leq\lambda\|Y_{k}-Z_{k}\|+\big{\|}\Psi_{h}(Z_{k+1}+\Psi_{-h}(Y_{k+1}))+\Psi_{-h}(Z_{k+1})\big{\|}+\|\Psi_{-h}\|_{\text{Lip-}1}\big{\|}Y_{k+1}-Z_{k+1}\big{\|}$
$\leq\lambda\|Y_{k}-Z_{k}\|+\big{\|}\Psi_{h}(Z_{k+1}+\Psi_{-h}(Z_{k+1}))+\Psi_{-h}(Z_{k+1})\big{\|}+C_{2}\,h\big{\|}Y_{k+1}-Z_{k+1}\big{\|}$
$\hskip 64.01869pt+\big{\|}\Psi_{h}(Z_{k+1}+\Psi_{-h}(Y_{k+1}))-\Psi_{h}(Z_{k+1}+\Psi_{-h}(Z_{k+1}))\big{\|}$
$\leq\lambda E_{k}^{YZ}+C_{3}h^{\alpha+1}+C_{2}hE_{k+1}^{YZ}+C_{2}^{2}h^{2}E_{k+1}^{YZ}\,.$
Therefore, we have the following inequality,
$E_{k+1}^{YZ}\leq\frac{1}{1-C_{2}h-C_{2}^{2}h^{2}}\big{(}\lambda E_{k}^{YZ}+C_{3}h^{\alpha+1}\big{)},$
provided $C_{2}h<\frac{1+\sqrt{5}}{2}$. Note that $(1-x-x^{2})^{-1}\leq 16^{x}$ for $x\in[0,\frac{1}{2}]$. Hence, for $h\leq\frac{1}{2C_{2}}$,
$E_{k+1}^{YZ}\leq 16^{C_{2}h}\big{(}\lambda E_{k}^{YZ}+C_{3}h^{\alpha+1}\big{)}.$
Applying the above inequality $k$ times and using $E_{0}^{YX}=0$, we have
$E_{k}^{YZ}$ $\leq 16^{C_{2}h}C_{3}\,h^{\alpha+1}\frac{16^{C_{2}kh}-1}{16^{C_{2}h}-1}$
$=C_{3}\frac{16^{C_{2}h}}{\log(16)C_{2}}\big{(}16^{C_{2}kh}-1\big{)}\frac{\log(16)C_{2}h}{e^{\log(16)C_{2}h}-1}h^{\alpha}$
$\leq\underbrace{C_{3}\frac{16^{C_{2}h}}{\log(16)C_{2}}\big{(}16^{C_{2}T}-1\big{)}}_{=:\,C_{4}}h^{\alpha},$
where we used $kh\leq T$ and the inequality $1+x\leq e^{x}$. ∎
Bringing the above analysis together, we now prove Theorem 3.1. The Theorem description is reproduced here with the notation used in this appendix.
Theorem.
For a fixed time horizon $T>0$, we consider the ODE (11) over the interval $[0,T]$. Let $T=Nh$ where $N>0$ denotes the number of steps and $h>0$ is the step size. Consider the reversible numerical solution $\{(Y_{k},Z_{k})\}_{0\leq k\leq N}$ with $Y_{0}=Z_{0}=y_{0}$ and, for $k\geq 0$,
$Y_{k+1}$ $:=\lambda Y_{k}+(1-\lambda)Z_{k}+\Psi_{h}(Z_{k}),$ (16)
$Z_{k+1}$ $:=Z_{k}-\Psi_{-h}(Y_{k+1}),$ (17)
where $\lambda\in(0,1]$ and $\Psi$ denotes an order $\alpha$ ODE solver satisfying the Lipschitz condition. Then there exist constants $h_{\max},C_{5}>0$ such that, for $h\in(0,h_{\max}]$,
$\|Y_{k}-y(t_{k})\|\leq C_{5}h^{\alpha}\,,$ (18)
for $0\leq k\leq N$ where $t_{k}:=kh$.
Proof.
Let $X_{k}:=\lambda^{N-k}Y_{k}+(1-\lambda^{N-k})Z_{k}$ and consider the error $E_{k}:=\|X_{k}-y(t_{k})\|$. Then
$E_{k+1}$ $=\big{\|}\lambda^{N-(k+1)}Y_{k+1}+(1-\lambda^{N-(k+1)})Z_{n+1}-y(t_{k+1})\big{\|}$
$=\big{\|}\lambda^{N-(k+1)}\big{(}\lambda Y_{k}+(1-\lambda)Z_{k}+\Psi_{h}(Z_{k})\big{)}+(1-\lambda^{N-(k+1)})\big{(}Z_{k}-\Psi_{-h}(Y_{k+1})\big{)}-y(t_{k+1})\big{\|}$
$=\big{\|}\lambda^{N-k}Y_{k}+(1-\lambda^{N-k})Z_{k}+\lambda^{N-(k+1)}\Psi_{h}(Z_{k})-(1-\lambda^{N-(k+1)})\big{(}\Psi_{-h}(Y_{k+1})\big{)}-y(t_{k+1})\big{\|}$
$\leq E_{k}+\big{\|}\,y(t_{k})+\lambda^{N-(k+1)}\Psi_{h}(Z_{k})-(1-\lambda^{N-(k+1)})\big{(}\Psi_{-h}(Y_{k+1})\big{)}-y(t_{k+1})\big{\|}$
$\leq E_{k}+\lambda^{N-(k+1)}\big{\|}\Psi_{h}(Z_{k})+\Psi_{-h}(Y_{k+1})\big{\|}+\big{\|}y(t_{k})-\big{(}y(t_{k+1})+\Psi_{-h}(Y_{k+1})\big{)}\big{\|}$
$\leq E_{k}+\lambda^{N-(k+1)}\big{\|}\Psi_{h}(Z_{k+1}+\Psi_{-h}(Y_{k+1}))+\Psi_{-h}(Y_{k+1})\big{\|}$
$\hskip 25.60747pt+\big{\|}y(t_{k})-\big{(}y(t_{k+1})+\Psi_{-h}(y(t_{k+1})\big{)}\big{\|}+\big{\|}\Psi_{-h}(y(t_{k+1}))-\Psi_{-h}(Y_{k+1})\big{\|}$
$\leq E_{k}+\lambda^{N-(k+1)}\big{\|}\Psi_{h}(Y_{k+1}+\Psi_{-h}(Y_{k+1}))+\Psi_{-h}(Y_{k+1})\big{\|}$
$\hskip 25.60747pt+\lambda^{N-(k+1)}\big{\|}\Psi_{h}(Z_{k+1}+\Psi_{-h}(Y_{k+1}))-\Psi_{h}(Y_{k+1}+\Psi_{-h}(Y_{k+1}))\big{\|}$
$\hskip 25.60747pt+C_{1}h^{\alpha+1}+\big{\|}\Psi_{-h}(y(t_{k+1}))-\Psi_{-h}(X_{k+1})\big{\|}+\big{\|}\Psi_{-h}(X_{k+1})-\Psi_{-h}(Y_{k+1})\big{\|}$
$\leq E_{k}+\lambda^{N-(k+1)}C_{3}h^{\alpha+1}+\lambda^{N-(k+1)}C_{2}h\|Z_{k+1}-Y_{k+1}\|$
$\hskip 25.60747pt+C_{1}h^{\alpha+1}+C_{2}h\,E_{k+1}+C_{2}h\|X_{k+1}-Y_{k+1}\|$
$\leq E_{k}+\lambda^{N-(k+1)}C_{3}h^{\alpha+1}+\lambda^{N-(k+1)}C_{2}h\|Y_{k+1}-Z_{n+1}\|$
$\hskip 25.60747pt+C_{1}h^{\alpha+1}+C_{2}h\,E_{k+1}+(1-\lambda^{N-(k+1)})C_{2}h\|Y_{k+1}-Z_{n+1}\|$
$\leq E_{k}+C_{2}h\,E_{k+1}+(C_{1}+C_{2}C_{4}+C_{3})h^{\alpha+1}.$
Therefore, provided $C_{2}h<1$,
$E_{k+1}\leq\frac{1}{1-C_{2}h}E_{k}+\frac{C_{1}+C_{2}C_{4}+C_{3}}{1-C_{2}h}h^{\alpha+1}.$
Since $(1-x)^{-1}\leq 1+2x\leq e^{2x}$ for $x\in[0,\frac{1}{2}]$, we have for $h\leq\frac{1}{2C_{2}}$,
$E_{k+1}\leq e^{2C_{2}h}E_{k}+(C_{1}+C_{2}C_{4}+C_{3})(1+2C_{2}h)h^{\alpha+1}.$
Applying the above inequality $k$ times and using $E_{0}=0$, we have
$E_{k}$ $\leq(C_{1}+C_{2}C_{4}+C_{3})(1+2C_{2}h)h^{\alpha+1}\frac{e^{2C_{2}kh}-1}{e^{2C_{2}h}-1}$
$=(C_{1}+C_{2}C_{4}+C_{3})(1+2C_{2}h)\frac{e^{2C_{2}kh}-1}{2C_{2}}\frac{2C_{2}h}{e^{2C_{2}h}-1}h^{\alpha}$
$\leq(C_{1}+C_{2}C_{4}+C_{3})(1+2C_{2}h_{\max})\frac{e^{2C_{2}T}-1}{2C_{2}}h^{\alpha},$
where we used $kh\leq T$ and the inequality $1+x\leq e^{x}$. Therefore
$\|Y_{k}-y(t_{k})\|$ $\leq\|Y_{k}-X_{k}\|+\|X_{k}-y(t_{k})\|$
$=E_{k}+(1-\lambda^{N-k})\|Y_{k}-Z_{k}\|$
$\leq\underbrace{\left((C_{1}+C_{2}C_{4}+C_{3})(1+2C_{2}h_{\max})\frac{e^{2C_{2}T}-1}{2C_{2}}+C_{4}\right)}_{=:\,C_{5}}h^{\alpha}.$
Appendix B Proof of stability (Theorem 3.2)
For convenience we provide the description of Theorem 3.2 here.
Theorem.
Let $\Psi$ be given by an explicit Runge-Kutta solver. Then the reversible numerical solution $\{y_{n},z_{n}\}_{n\geq 0}$ given by
$y_{n+1}$ $=\lambda y_{n}+(1-\lambda)z_{n}+\Psi_{h}(t_{n},z_{n}),$
$z_{n+1}$ $=z_{n}-\Psi_{-h}(t_{n+1},y_{n+1}),$
is linearly stable iff
$|\Gamma|<1+\lambda,$ (19)
where
$\Gamma=1+\lambda-(1-\lambda)R(-h\alpha)-R(-h\alpha)R(h\alpha).$ (20)
Proof.
Firstly, we re-write the reversible method applied to the linear test problem (Definition 3.1) using the Runge-Kutta transfer function. We get
$y_{n+1}$ $=\lambda y_{n}+(1-\lambda)z_{n}+R(h\alpha)z_{n},$ (21)
$z_{n+1}$ $=z_{n}-R(-h\alpha)y_{n+1}.$
Writing the coupled system as a matrix-vector product results in
$\begin{pmatrix}y_{n+1}\\ z_{n+1}\end{pmatrix}=\begin{pmatrix}\lambda&1-\lambda+R(h\alpha)\\ -\lambda R(-h\alpha)&1+(1-\lambda)R(h\alpha)-R(-h\alpha)R(h\alpha)\end{pmatrix}\begin{pmatrix}y_{n}\\ z_{n}\end{pmatrix}.$ (22)
For stability we require that the eigenvalues of the coupling matrix are by magnitude less than one. The eigenvalues are given by the characteristic equation,
$P(e_{k})=e_{k}^{2}-\left(1+\lambda-(1-\lambda)R(-h\alpha)-R(-h\alpha)R(h\alpha)\right)e_{k}+\lambda=0.$ (23)
Routh-Hurwitz test
The Routh-Hurwitz test is a condition on the coefficients of a polynomial such that the roots lie in the left half complex plane. Instead we require that the roots lie in the unit disc. Therefore, we apply the following map to the characteristic equation $P(e_{k})$,
$Q(e_{k})=(e_{k}-1)^{2}P\left(\frac{e_{k}+1}{e_{k}-1}\right),$ (24)
such that if $Q$ is Routh-Hurwitz stable, then the roots of $P$ lie in the open unit disc (referred to as Schur stable) [20].
The polynomial $Q$ obtained by the transformation is given by
$Q(e_{k})=\left(1+\lambda-\Gamma\right)e_{k}^{2}+2(1-\lambda)e_{k}+(1+\lambda+\Gamma)=0,$ (25)
where
$\Gamma=1+\lambda-(1-\lambda)R(-h\alpha)-R(-h\alpha)R(h\alpha).$ (26)
We proceed to check when $Q$ is Routh-Hurwitz stable. For quadratic polynomials, the Routh-Hurwitz test reduces to checking that the coefficients of $Q$ are all positive. Therefore, for stability, we require
[…]
Condition (ii) is immediately satisfied by the definition of $\lambda\in(0,1)$. Condition (i) and (iii) may be combined into the condition
$|\Gamma|<1+\lambda.$ (28)
Appendix C Experimental details
All experiments were run on one Dual NVIDIA RTX A6000, 48 GB.
C.1 Toy Problem
Regression task
Regression was performed on the cubic polynomial function, $y=f(x)=x^{3}+x$. The dataset $D=\{x_{i},y_{i}\}_{0\leq i\leq n}$ was generated with $n=1000$.
Neural ODE
The vector field of the Neural ODE was given by a simple feed-forward network with initial linear layer, $\tanh$ activation and final linear layer. The width of the linear layers was $50$.
Gradient calculation
The Neural ODE gradients were calculated using the reversible method, optimise-then-discretise and discretise-then-optimise. In each method the Midpoint scheme was used; the step size was $h=0.01$ and the integration time was varied from $T=1.0$ to $5.0$. The coupling parameter for the reversible method was set to $\lambda=0.99$.
For each method gradients were calculated for $10$ runs at each integration time $T$. The $L^{1}$ gradient errors and memory usages in Figure 2 at each integration time $T$ are the means taken over the $10$ runs.
C.2 Continuous Normalising Flows
Dataset
The dataset used for the experiments in Table 1 was the two moons dataset. This can be imported from the sklearn Python library [22]. The number of points generated was $n=1000$.
Neural ODE
The vector field of the Neural ODE was given by a feed-forward network with $3$ ConcatSquash layers (introduced in [21]), each followed by a $\tanh$ activation function, and $1$ final ConcatSquash layer. The width of the ConcatSquash layers was $128$.
Training
The Neural ODE was trained using the reversible method and discretise-then-optimise for comparison. For both training methods, a selection of numerical solvers were used: Midpoint, Runge-Kutta 4 (RK4) and Dormand-Prince 5/4 (Dopri5) [28]. In every case, the step size $h=0.01$ and the integration time $T=1.0$. The coupling parameter for the reversible method was set to $\lambda=0.999$. Training proceeded for $5000$ steps using the AdamW optimiser with learning rate $10^{-3}$ and weight decay parameter $10^{-5}$. For each method and numerical solver, training was repeated $10$ times with random network initialisations.
C.3 Neural CDE
Dataset
The CharacterTrajectories dataset was used for the Neural CDE experiment, consisting of the $(x,y)$ velocity trajectory and pen tip force while writing a Latin alphabet character. The maximum time-series (trajectory) length was $182$ and all time-series were zero-padded to this length. The training set size was $2240$ time-series examples and the test set size was $618$ examples.
Control path
Each time-series was interpolated using Backward Hermite Cubic Splines [29], with $182$ knots, to form a control path $X(t)$.
Neural CDE
The vector field of the Neural CDE was a Multi-Layer Perceptron (MLP) with hidden size $32$, width size $32$ and depth $4$. The latent state (size $32$) of the Neural CDE was initialised by a learnt linear transformation of the initial input path value $X(0)$.
Training
The Neural CDE was trained using the reversible method and discretise-then-optimise for comparison. For both training methods, a selection of numerical solvers were used: Midpoint, Runge-Kutta 4 (RK4) and Dormand-Prince 5/4 (Dopri5) [28]. In every case, the step size $h=0.01$ and the integration time $T=1.0$. The coupling parameter for the reversible method was set to $\lambda=0.999$. Training proceeded for $1000$ steps using the AdamW optimiser with learning rate $10^{-3}$ and weight decay parameter $10^{-5}$. For each method and numerical solver, training was repeated $10$ times with random network initialisations.